in a billiard table a ball of mass weighing 250 g rolls at a velocity of 2?5 m/s and rebounds back at 2.5 m/s what is it change in momentum

1
Expert's answer
2021-11-09T16:36:04-0500

We can find the change in momentum as follows:


"\Delta p =p_f-p_i=mv_f-mv_i=m(v_f-v_i),""\Delta p =0.25\ kg\times(-2.5\ \dfrac{m}{s}-2.5\ \dfrac{m}{s})=-1.25\ \dfrac{kg\times m}{s}."